Grade 10Physical Chemistry

One litre of oxygen at STP made to react with three times the volume of carbon monoxide at STP.Find the composition of the resultant mixture.

Piyali Kundu
10 Years agoGrade 10

2 Answers

Vikas TU
9 Years ago
CO              +              1/2O2             ----------------->             CO2
3L                                  1L                                        
From mole analysis,
O2 is L.R.
Therfore,
Volume of CO2 is1/0.5 = 2 Litre.
Misba
4 Years ago
1CO+1/2 O2    => CO2
1 mole CO.       => 0.5 mole O2
3 litre CO.         => 3×0.5 = 1.5 litre of O2
   .         
.     .   Limiting reagent is O2 
0.5 O2.       =>  1 litre CO 
1 litre O2.   =>  1×1÷1/2 = 2 litre
n = 1/22.4 = W/28
W = 1.25 gram